1/x + 1/y = 5, can you solve for x without ending up with an xy
Found 2 solutions by jim_thompson5910, nerdybill:
Answer by jim_thompson5910(35256)
Answer by nerdybill(7384)

Not sure what you mean but here's how to solve for x:
1/x + 1/y = 5
multiply both sides by xy:
(xy)(1/x + 1/y) = 5xy
distribute on the left:
y + x = 5xy
y = 5xy - x
factor the right:
y = x(5y - 1)
y/(5y-1) = x